# Deciding Subspace reachability problems with application to Skolem’s Problem

### v1

@@ Abstract.
The higher-dimensional version of Kannan and Lipton’s Orbit Problem asks whether it is decidable if a target vector space can be reached from a starting point under repeated application of a linear transformation. This problem has remained open since its formulation, and in fact generalizes Skolem’s Problem — a long-standing open problem concerning the existence of zeros in linear recurrence sequences. Both problems have traditionally been studied using algebraic and number theoretic machinery. In contrast, this paper reduces the Orbit Problem to an equivalent version in real projective space, introducing a basic geometric reference for examining and deciding problem instances. We find this geometric toolkit enables basic proofs of sweeping assertions concerning the decidability of certain problem classes, including results where the only other known proofs rely on sophisticated number-theoretic arguments.
@@
The author is supported by the National Science Foundation Graduate Research Fellowship Program under Grant No. 2140001.
@@ 1. Introduction
In a pair of seminal papers [kannan1980orbit, kannan1986polynomial] Kannan and Lipton introduced the Orbit Problem, motivated by reachability questions of linear sequential machines raised by Harrison in 1969 [harrison1969lectures]. Given a linear transformation $A\in\mathbb{Q}^{d\times d}$, and elements $x,y\in\mathbb{Q}^{d}$, the Orbit Problem asks whether it can be decided if there exists an $n\in\mathbb{N}$, such that $A^{n}x=y$. Kannan and Lipton proved decidability of this point-to-point version of the problem, but remarked that when the target is a subspace $U$ of $\mathbb{Q}^{d}$, the problem becomes far more difficult. Despite the importance of this problem due to its connection with Skolem’s Problem and areas of program verification, no progress was made on the higher-dimensional Orbit Problem until 2013, when Chonev, Ouaknine, and Worrell proved that the higher-dimensional Orbit Problem is decidable whenever the dimension of the target space is one, two, or three [chonev2013orbit, chonev2016complexity]. In following work, the same authors considered a version of the problem where the target space is not a subspace but rather a polytope [chonev2014polyhedron], proving decidability when the target is of dimension three or less, and hardness with respect to long-standing number theoretic open problems for higher dimensions.
Since this set of breakthroughs however, little progress has been made on the higher-dimensional Orbit Problem, due in large part to a severe lack of structure. There have been results establishing decidability for generalized cases in which the source and target sets are both either polytopes [almagorPolytopeCollision] or semi-algebraic sets [almagorSemialgebraicOrbit], however, such results continue to only apply to the case when the dimension is at most three. An alternative approach present in recent literature involves generating sets invariant under the action of the matrix $A$ that contain the starting point, and then proving such sets are disjoint from the target set [de2018left, fijalkow2019complete, almagor2022minimal]. However, the non-existence of an invariant set does not imply decidability, and these results either only apply to instances for which decidability is already known, or fail to capture natural classes of problem instances and are conditioned on long-standing number theoretic conjectures. This body of literature constitutes the state-of-the-art, speaking to the difficulty of the problem.
Our work advances the state-of-the-art by providing a geometric machinery used to determine a broad category of decidable instances of the higher-dimensional Orbit Problem, not by restricting the dimension of the target and ambient spaces, but rather the spectral structure of the matrix. The main result of this paper, stated below, concerns a generalization of the higher-dimensional Orbit Problem where the target space is a union of subspaces with convex polytopes.
@@
Let $(A,x,U)$ be any non-degenerate [^1] An instance is non-degenerate if the matrix $A$ has no two distinct eigenvalues whose quotient is a root of unity. It is well-known that any degenerate instance can be effectively reduced to solving a collection of non-degenerate instances. instance of the Orbit Problem, where $A\in\mathbb{Q}^{d\times d}$, $x\in\mathbb{Q}^{d}$ is non-trivial [^2] We say a point $x$ is non-trivial (w.r.t. a matrix $A$) if $x$ has a non-zero component with respect to the Jordan blocks of Jordan matrix $J(A)$. The non-triviality condition on $x$ assures the problem does not collapse to a lower-dimension; if the condition is removed an essentially identical statement holds., and $U$ is a target set in $\mathbb{Q}^{d}$ composed of a finite union of subspaces with convex polytopes, described by rational parameters.
Suppose $A$ has $r$ distinct eigenvalues $\lambda_{1},...,\lambda_{r}$ of maximal modulus. Let $n_{i}$ denote the algebraic multiplicity of the root $\lambda_{i}$, $i=1,...,r$, in the minimal polynomial of $A$. Suppose $n_{1}=\cdots=n_{p}$, $p\leq r$, are the largest such values. Then there exists a subspace $W$ computable from $A$ and $x$ of dimension $p$, such that if $W\cap U=\{0\}$, it is decidable whether there exists $n\in\mathbb{N}$ such that $A^{n}x\in U$.
Intuitively, Theorem 1 can be thought of as solving “typical” instances of the Orbit Problem whenever $U$ is a subspace and $p+\text{dim}(U)\leq d$. But the Theorem is to be primarily viewed as providing a tool for determining decidability of problem classes. For instance, one immediate application of Theorem 1 is in proving the result of Chonev, Ouaknine, and Worrell that the higher-dimensional Orbit Problem is decidable whenever the dimension of the target space is one [chonev2013orbit, chonev2016complexity]. Namely, we use Theorem 1 to give a simple geometric proof of the following
@@
Let $(A,x,U)$ be any non-degenerate instance of the Orbit Problem, where $A\in\mathbb{Q}^{d\times d}$ is nonsingular, $x\in\mathbb{Q}^{d}$, and $U=\emph{span}(v)$ for $v\in\mathbb{Q}^{d}\setminus\{0\}$. Then it is decidable whether there exists an $n\in\mathbb{N}$ such that $A^{n}x\in U$.
The use of Theorem 1 reaches beyond the Orbit Problem to Skolem’s Problem as well, which is known to reduce to the higher-dimensional Orbit Problem. Skolem’s Problem is a long-standing open question asking whether it is decidable if the set of zeros of a linear recurrence sequence is empty (see e.g. [ouaknine2015linear, halava2005skolem] for review), and is known to be an impenetrable problem where even simple cases escape our most advanced techniques. Indeed, Tao described the openness of this problem as “faintly outrageous,” as it indicates “we do not know how to decide the halting problem even for linear automata” [tao2008structure].
Decidability of Skolem’s Problem for linear recurrence sequences of dimension three and four was given in the 1980’s [shorey1984distance, vereshchagin1985occurrence], along with decidability when there are at most three dominating characteristic roots of the sequence (see [sha2019effective] for refinements of this work). The methods used to prove such results rely crucially on sophisticated results in transcendental number theory, particularly Baker’s lower bounds on the magnitudes of linear forms in logarithms of algebraic numbers, and van der Poorten’s results in the setting of $p$-adic valuations. In contrast, we use Theorem 1 paired with a basic geometric insight to prove decidability of Skolem’s Problem for the case of arbitrarily many dominating roots, so long as the set of dominating roots contains a real root with largest algebraic multiplicity in the minimal polynomial.
@@
Let $\{u_{n}\}_{n=0}^{\infty}$ be an order $d$ non-degenerate linear recurrence sequence, with $A$ the associated matrix, and suppose the initial terms form a vector $x\in\mathbb{Q}^{d}$ non-trivial with respect to the matrix $A$. Suppose the linear recurrence sequence has $r\leq d$ distinct characteristic roots $\lambda_{1},...,\lambda_{r}$ of largest modulus. Let $n_{i}$ denote the algebraic multiplicity of the root $\lambda_{i}$, $i=1,...,r$, in the minimal polynomial of $A$. Suppose $\lambda_{1}$ is real, and $n_{1}>n_{i}$, $i=2,...,r$. Then it is decidable whether the sequence has a zero term.
We emphasize that in both Theorem 1 and 3, we care only about the multiplicity of the roots in the minimal polynomial of $A$, which need not coincide with the multiplicity of the roots of the characteristic polynomial. Moreover, in contrast to many of the results towards Skolem’s problem such as [sha2019effective, biluMFCSskolem], we highlight that Theorem 3 makes no assumptions about the simplicity (diagonalizability) of the linear recurrence sequence, nor about the order of the sequence.
1.1. Techniques
@@
Traditionally, work toward determining the decidability of the Orbit Problem and related problems such as Skolem’s Problem has leveraged heavy machinery from various domains of algebraic geometry, mathematical logic, and number theory — particularly algebraic and transcendental number theory. In contrast, our approach is to develop a machinery suitable for application to the Orbit Problem by leveraging basic geometric and dynamical insight.
@@
In essence, we consider not the dynamical system $(\mathbb{R}^{d},A)$ the Orbit Problem is formulated in, but rather the associated induced dynamical system over real projective space (i.e. after projecting onto $S^{d-1}$ and identifying opposite points). Working instead in $\mathbb{R}\mathbb{P}^{d-1}$, we can ask an equivalent version of the Orbit Problem that is decidable if and only if the original formulation is decidable. We find that by compactifying the state space in such a way, the asymptotic behavior of the system becomes far more understandable: in projective space, we have asymptotically stable (converging) orbits, while no-such behavior is expressed in Euclidean space. As a consequence, we can use arguments unavailable in the original formulation. The central contribution of our paper can then be interpreted as a fine-grained examination of the $\omega$-limit sets — or more generally the stable manifolds — of dynamical systems $(\mathbb{R}\mathbb{P}^{d-1},f)$ for $f\in PGL(\mathbb{R}^{d})$, namely linear dynamical systems over real projective space, with this deepened understanding lending itself to deciding the higher-dimensional Orbit Problem.
@@
This reduction is both surprising and noteworthy for a number of reasons. First, we note that although we lose an entire manifold dimension after reducing the problem to projective space, we maintain all the necessary information for deciding the Orbit Problem — indeed, the reduction can be seen as keeping only the “essential” aspects of the Orbit Problem. In addition, this reduction compactifies the state space, and as a consequence enables the application of general tools and intuition from the study of continuous dynamical systems over a compact space, for which there is far more structure and results than in the unbounded case, enabling the application of an entire area of mathematics not used to date.
@@ 1.2. Related Work
One of the original sources of motivation for studying the Orbit Problem was Skolem’s Problem, due to the fact that Skolem’s Problem reduces to the Orbit Problem when the target space has dimension $d-1$. Although Skolem’s Problem is easy to describe, it is difficulty to prove deep theorems on account of a lack of machinery. Indeed, in terms of lower bounds, it is known that Skolem’s Problem is NP-hard [blondel2002presence] — which translates to the Orbit Problem when restriction is not placed on the dimension of the target subspace. Moreover, deep work has shown that further advances on the Skolem and Orbit Problem for higher dimensions are likely to be hard due to the fact that such advances in related problems would entail major advances in Diophantine approximation [ouaknine2014positivity]. Despite these difficulties, there continues to be notable advances toward Skolem’s problem (see e.g. [lipton2022skolem, biluMFCSskolem]), including recent promising new techniques by way of Universal Skolem Sets [kenison2020skolem, luca2021universal, lucaMFCSuniversal, luca2023skolem]. Nonetheless, despite the intense study on this problem, our understanding remains incomplete.
@@ Acknowledgements
The author is grateful to David Cash for the continued feedback and support.
2. Preliminaries
The purpose of this section is to establish notation, as well as results and machinery used for the proofs of Theorems 1, and 3 in Section 3. Sections 2.3 and 2.4 are the richest, where the key lemmas of this paper are formed.
Given an instance $(A,x,U)$ of the Orbit Problem, always take $A\in\mathbb{Q}^{d\times d}$, $x\in\mathbb{Q}^{d}$, and $U$ presented via sets of basis vectors from $\mathbb{Q}^{d}$. With such a basis in hand it is clearly decidable whether a vector $A^{n}x\in\mathbb{Q}^{d}$ is in $U$. For instance, to verify whether $A^{n}x\in U$ when $U$ is just a subspace, we compute the determinant $B^{T}B$ where $B$ is the matrix whose columns are $A^{n}x$ and the basis vectors specifying $U$. Then $n$ is a witness to the problem if and only if this determinant is zero.

### v2

@@ Abstract.
The higher-dimensional version of Kannan and Lipton’s Orbit Problem asks whether it is decidable if a target subspace can be reached from a starting point under repeated application of a linear transformation. Similarly, the continuous analog of the Orbit Problem asks if a flow induced by a linear system of differential equations ever reaches some specified subspace. The decidability of both problems remains open, and in fact the problems generalize the discrete and continuous versions of Skolem’s Problem. The object of this paper is to communicate a geometric perspective of the discrete and continuous Orbit Problems, alternate to the traditional and highly technical algebraic and number-theoretic approaches to the problem. We derive a simple decision procedure capable of deciding a certain class of instances of the Orbit Problem, and, as an application, we obtain alternate proofs to a number of results using elementary geometric arguments.
@@ 1.1. Orbit Problem Overview
In a pair of seminal papers [kannan1980orbit, kannan1986polynomial] Kannan and Lipton introduced the Orbit Problem, motivated by reachability questions of linear sequential machines raised by Harrison [harrison1969lectures]. Given a linear transformation $A\in\mathbb{Q}^{d\times d}$, and elements $x,y\in\mathbb{Q}^{d}$, the Orbit Problem asks whether it can be decided if there exists an $n\in\mathbb{N}$, such that $A^{n}x=y$. Kannan and Lipton proved decidability of this point-to-point version of the problem, but remarked that when the target is a subspace $U$ of $\mathbb{Q}^{d}$, the problem becomes far more difficult. One may additionally consider a continuous version of the Orbit Problem, in which it is to be decided whether a flow induced by a linear system of differential equations reaches a target hyperplane, or whether real-valued exponential polynomials have any real roots [bell2010continuous]. More formally, presented with an ordinary differential equation $dx(t)/dt=Ax(t)$, with $A\in\mathbb{Q}^{d\times d}$, and initial condition $x(0)\in\mathbb{Q}^{d}$, it is to be decided whether there exists a $t\in\mathbb{R}^{+}$ such that $c^{\top}\exp(At)x(0)=0$ for $c\in\mathbb{Q}^{d}$. Despite the importance of these problems due to their connection with Skolem’s Problem and the foundations of program verification and control theory, progress on the problems has been slow.
In the discrete-time setting, no progress was made on the higher-dimensional Orbit Problem until 2013, when in beautiful work Chonev, Ouaknine, and Worrell proved that the higher-dimensional Orbit Problem is decidable whenever the dimension of the target space is one, two, or three [chonev2013orbit, chonev2016complexity]. In following work, the same authors considered a version of the problem where the target space is not a subspace but rather a polytope [chonev2014polyhedron], proving decidability when the target is of dimension three or less, and hardness with respect to long-standing number theoretic open problems for higher dimensions. In related work, decidability for dimensions up to three was also shown for generalized cases in which the source and target sets are both either polytopes [almagorPolytopeCollision] or semi-algebraic sets [almagorSemialgebraicOrbit]. One other prominent recent approach involves generating sets that are invariant under the action of the matrix $A$ that contain the starting point, and then proving such sets are disjoint from the target set [de2018left, fijalkow2019complete, almagor2022minimal]. However, many difficulties in analyzing the Orbit Problem remain very-much alive, with most substantive progress being made conditioned on long-standing number theoretic conjectures. In particular, the decision procedures provided in these works rely on highly technical methods from number theory, model theory, and algebra.
The continuous-time setting offers no reprieve. In fact there have been fewer results making progress on the continuous version of the Orbit problem; partially because the usual algebraic and number theoretic techniques do not carry over to the continuous time setting as easily. Nevertheless, there have been a number of strong results toward decidability. Early work of Hainry [hainry2008reachability, hainry2009decidability] proved decidability of the continuous point-to-point Orbit Problem, a result which has since been strengthened [dantam2021decidability, chen2015continuous]. Bell et al. [bell2010continuous] made substantial progress on the problem, particularly proving NP-hardness and decidability for dimensions two or less. In another successful line of work, Chonev, Ouaknine, and Worrell proved decidability up to dimension eight conditioned on Schanuel’s Conjecture, and proved decidability up to dimension eight unconditionally for a restricted version of the problem [chonev2016recurrent, chonev_et_al2016, 10.1145/3603543]. Despite these deep results toward the problem, there remain many obstacles to further progress.
1.2. Statement of Results
The aim of this paper is to present an approach to the discrete and continuous Orbit Problems centered around the observation that equivalent problems can be asked and answered in real-projective space rather than Euclidean space. Once this observation has been made, many technical theorems become relatively elementary to prove due to the substantial increase in geometric structure available.
This paper begins development in the discrete-time setting, with first application of our techniques to the discrete-time Orbit Problem and Skolem Problem. Subsequent application of our techniques in the context of flows is straight-forward. Our first result is concerned with developing a simple geometric algorithm capable of deciding the problem when certain spectral and geometric constraints are satisfied. However, the following result is principally used as a foothold in proving all the other results in this paper.
@@ 1.2. Statement of Results
Let $(A,x,U)$ be any instance of the discrete-time Orbit Problem, where $A\in\mathbb{Q}^{d\times d}$, $x\in\mathbb{Q}^{d}$ is non-trivial [^1] We say a point $x$ is non-trivial (w.r.t. a matrix $A$) if $x$ has a non-zero component with respect to the Jordan blocks of Jordan matrix $J(A)$. The non-triviality condition on $x$ assures the problem does not collapse to a lower-dimension; if the condition is removed an essentially identical statement holds., and $U$ is a target subspace in $\mathbb{Q}^{d}$, described by rational parameters. Suppose $A$ has $r$ distinct eigenvalues $\lambda_{1},\dots,\lambda_{r}$ of largest modulus. Let $n_{i}$ denote the multiplicity of the root $\lambda_{i}$, $i=1,\dots,r$, in the minimal polynomial of $A$. Suppose $n_{1}=\cdots=n_{p}$, $p\leq r$, are the largest such values. Then there exists a subspace $W$ computable from $A$ and $x$ of dimension $p$, such that if $W\cap U=\{0\}$, it is decidable whether there exists $n\in\mathbb{N}$ such that $A^{n}x\in U$.
Once the step of compactifying state space has been taken by translating the problem to projective space, proof of Theorem 1 is conceptually easy: it essentially reduces to a case analysis of the asymptotics of any orbit under iteration of a matrix in Jordan normal form. In particular, Jordan blocks comprising the normal form of the matrix $A$ carry all the information needed to determine the subspace any given initial condition will develop toward, after compactifying by projecting onto real projective space.
Theorem 1 is to be primarily viewed as providing a tool for determining decidability of problem classes. For instance, one immediate application of Theorem 1 is in proving a weakened version of the breakthrough result of Chonev, Ouaknine, and Worrell that the higher-dimensional Orbit Problem is decidable whenever the dimension of the target space is one [chonev2013orbit, chonev2016complexity]. Namely, we use Theorem 1 to give a proof of the following theorem using only elementary geometric arguments.
@@
Let $(A,x,U)$ be any instance of the Orbit Problem such that all eigenvalues of $A\in\mathbb{Q}^{d\times d}$ have distinct modulus up to complex conjugate pairs, $x\in\mathbb{Q}^{d}$, and $U$ a subspace of dimension at most one. Then it is decidable whether there exists an $n\in\mathbb{N}$ such that $A^{n}x\in U$.
The use of Theorem 1 reaches beyond the discrete-time Orbit Problem to Skolem’s Problem as well (see e.g. [ouaknine2015linear, halava2005skolem] for review), which is known to be an impenetrable problem where even simple cases escape our most advanced techniques. Indeed, Tao described the openness of this problem as “faintly outrageous,” as it indicates “we do not know how to decide the halting problem even for linear automata” [tao2008structure].
Decidability of Skolem’s Problem for linear recurrence sequences of dimension three and four was given in the 1980’s [shorey1984distance, vereshchagin1985occurrence], along with decidability when there are at most three dominating characteristic roots of the sequence (see [sha2019effective] for refinements of this work). The methods used to prove such results rely crucially on sophisticated results in transcendental number theory, particularly Baker’s lower bounds on the magnitudes of linear forms in logarithms of algebraic numbers, and van der Poorten’s results in the setting of $p$-adic valuations. In contrast, we use Theorem 1 paired with a basic geometric insight to provide an alternate proof of decidability of Skolem’s Problem for the case of arbitrarily many roots of largest modulus, so long as the set of such dominating roots contains a real root with largest multiplicity in the minimal polynomial.
@@
Let $\{u_{n}\}_{n=0}^{\infty}$ be an order $d$ non-degenerate linear recurrence sequence, with $A$ the companion matrix, and suppose the initial terms form a vector $x\in\mathbb{Q}^{d}$ non-trivial with respect to the matrix $A$. Suppose the linear recurrence sequence has $r\leq d$ distinct characteristic roots $\lambda_{1},\dots,\lambda_{r}$ of largest modulus. Let $n_{i}$ denote the multiplicity of the root $\lambda_{i}$, $i=1,\dots,r$, in the minimal polynomial of $A$. Suppose $\lambda_{1}$ is real, and $n_{1}>n_{i}$, $i=2,\dots,r$. Then it is decidable whether the sequence has a zero term.
We find the continuous version of the Orbit Problem to be amenable to our techniques as well. This is because real projective space produces a rather forgiving environment for analyzing flows, and particularly for bounding flows into certain regions of space. To this end, we give an elementary proof of decidability of the state of the art (unconditional) two-dimensional continuous-time Orbit Problem (Continuous Skolem-Pisot Problem) established in [bell2010continuous]—in fact we prove a strengthened version of the statement given in [bell2010continuous].
Theorem 4.
Let $dx(t)/dt=Ax(t)$, $A\in Gl(2,\mathbb{A})$, be an autonomous linear differential equation with $x(0)\in\mathbb{A}^{2}$, and let $U\subset\mathbb{A}^{2}$ be a line. If the eigenvalues of $A$ are real, it is decidable if the flow intersects $U$. If the eigenvalues of $A$ are complex, the flow intersects $U$ infinitely often.
We move to describe our techniques, in particular the basic geometric insight behind the results in this paper.
1.3. Overview of Techniques
@@
Traditionally, work toward the Orbit Problem has leveraged heavy machinery from various domains of algebraic geometry, mathematical logic, and number theory—particularly algebraic and transcendental number theory. In contrast, our approach is to develop a machinery suitable for application to the Orbit Problem by leveraging basic geometric and dynamical insight.
In essence, we consider not the dynamical system $(\mathbb{R}^{d},A)$ the Orbit Problem is formulated in, but rather the associated induced dynamical system over real projective space (i.e. after projecting onto $S^{d-1}$ and identifying opposite points), with the same compactification performed in the continuous-time analog. Working instead in $\mathbb{R}\mathbb{P}^{d-1}$, we can ask equivalent versions of the Orbit Problem that are decidable if and only if the original formulation is decidable.
@@
We find that by compactifying the state space in such a way, the asymptotic behavior of the system becomes far more understandable: in projective space, we have asymptotically stable (converging) orbits, while no-such behavior is expressed in Euclidean space. As a consequence, we can use arguments unavailable in the original formulation. The central contribution of our paper can then be interpreted as a fine-grained examination of the $\omega$-limit sets—or more generally the stable manifolds—of dynamical systems $(\mathbb{R}\mathbb{P}^{d-1},f)$ for $f\in PGL(\mathbb{R}^{d})$, namely linear maps and flows over real projective space, with this deepened understanding lending itself to deciding the discrete and continuous-time Orbit Problems. In practice, working directly with real-projective space is not feasible, so instead we consider the evolution of the angles between an orbit of a point and some subspace, which has the same desired effect. To make proofs fully rigorous and to speak of “convergence” of orbits, we then translate between consideration of angles between a point and subspace, and the convergence of a sequence in projective space to a cycle or fixed point.
@@
This reduction to projective space is both surprising and noteworthy for a number of reasons. First, we note that although we lose an entire manifold dimension after compactifying the space, we maintain all the necessary information for deciding the Orbit Problem—indeed, the reduction can be seen as keeping only the “essential” aspects of the Orbit Problem. In addition, the compactification enables the application of general tools and intuition from the study of continuous dynamical systems over a compact space, for which there is far more structure and results than in the unbounded case, enabling the application of an entire area of mathematics not used to date.
In the setting of flows, asymptotic behavior of autonomous linear differential equations is dictated by the Lyapunov exponents. These exponents determine asymptotic behavior of flows by the magnitudes and signs of the real parts of the eigenvalues. When projecting flows in $\mathbb{R}^{d}$ onto $\mathbb{RP}^{d-1}$, all flows are asymptotic to stationary points or cycles, since the compact phase space $\mathbb{RP}^{d-1}$ precludes divergence of trajectories. If a flow passes from one side to another of a target hyperplane then it must intersect it. By using this fact one may obtain stronger results than in the case of maps where asymptotic behaviors, even in real projective space, are far more subtle. We abuse this additional geometric structure present in the study of flows in $\mathbb{RP}^{d-1}$ when proving Theorem 4.
@@ 1.4. Related Work
One of the original sources of motivation for studying the Orbit Problem was Skolem’s Problem, due to the fact that Skolem’s Problem is a special case of the Orbit Problem. Although Skolem’s Problem is easy to describe, it is difficulty to prove deep theorems on account of a lack of machinery. Indeed, in terms of lower bounds, it is known that Skolem’s Problem is NP-hard [blondel2002presence]. This translates to the Orbit Problem when restriction is not placed on the dimension of the target subspace. Moreover, deep work has shown that further advances on the Skolem and Orbit Problem for higher dimensions are likely to be hard due to the fact that such advances in related problems would entail major advances in Diophantine approximation [ouaknine2014positivity]. Despite these difficulties, there continues to be notable advances toward Skolem’s problem (see e.g. [lipton2022skolem, biluMFCSskolem]), including recent promising new techniques by way of Universal Skolem Sets [kenison2020skolem, luca2021universal, lucaMFCSuniversal, luca2023skolem].
@@ Acknowledgements
The author is grateful to David Cash for the continued feedback and support. The author is supported by the National Science Foundation Graduate Research Fellowship Program under Grant No. 2140001.
2. Preliminary results
The purpose of this section is to establish notation, as well as results and machinery used for the proofs of Theorems 1,3, and 4 in Section 3. Sections 2.3 and 2.4 are the richest, where the key lemmas of this paper are formed. Specifically, we focus on development in the discrete-time setting for the discrete Orbit Problem. Once the results of this setting are established, techniques and intuition translate to the continuous setting in a very straight-forward way.
Given an instance $(A,x,U)$ of the Orbit Problem, always take $A\in\mathbb{Q}^{d\times d}$, $x\in\mathbb{Q}^{d}$, and $U$ presented via sets of basis vectors from $\mathbb{Q}^{d}$. With such a basis in hand it is clearly decidable whether a vector $A^{n}x\in\mathbb{Q}^{d}$ is in $U$.